OpenStudy (anonymous):

The weight W of an object varies inversely as the square of the distance d from the center of the earth. At sea level (3978 mi from the center of the earth), an astronaut weights 122 lb. Find her weight when she is 186 mi above the surface of the earth and the spacecraft is not in motion. Her weight is _ lb

OpenStudy (amistre64):

so we are using 3978 as our earth center distance eh... some books use 4000

OpenStudy (amistre64):

inversely proportional means that as one gets more, the other gets less. and we solve it with a fraction: y = 1/x is an inverse proportion

OpenStudy (amistre64):

we also use a number "k" as a constant in the equation to even the two out, otherwise itsd just be nonesense

OpenStudy (anonymous):

............this is deep

OpenStudy (amistre64):

W = k/(d^2) 122 = k/(3978)^2 k = 122((3978)^2) k = 19,305,870,48

OpenStudy (amistre64):

lol.... its in space, thats like the opposite of deep :)

OpenStudy (amistre64):

W = 19,305,870,48 // d^2 W = 19,305,870,48 // (3978+186)^2 W = 19,305,870,48 // 17,338,896 W = 111.344 lbs if I did it right

OpenStudy (anonymous):

.........

OpenStudy (anonymous):

YES!

OpenStudy (amistre64):

is that a right or a wrong lol if its the same program as the other girl that was here; try 111.34

OpenStudy (amistre64):

i take that for a correct lol

OpenStudy (anonymous):

Can you that one more time?

OpenStudy (anonymous):

My friend Kelley is on here.......and has the same question, just different numbers

OpenStudy (amistre64):

I dont know..... the rules change with every iteration of doing the math so it could get tricky.......... ;)

OpenStudy (anonymous):

Lol, she is pissed that you got it right for me

OpenStudy (amistre64):

hah!! its these big brains in this tiny little head lol

OpenStudy (anonymous):

Hey, it seems to work.........ok hers is.....

OpenStudy (anonymous):

her astronaut weights 130 lbs

OpenStudy (anonymous):

this stuff is nuts

OpenStudy (amistre64):

if I start throwing "pi" around like a clown, just shoot me!!

OpenStudy (anonymous):

lol ........wow....... math joke.

OpenStudy (amistre64):

is that the only difference in numbers? 130 lbs?

OpenStudy (anonymous):

One more thing her astronaut is 173 miles above the surface of the earth

OpenStudy (amistre64):

ok.... let me crunch some numbers, do the hokey pokey and turn myself about.....

OpenStudy (anonymous):

What is the hokey pokey really IS what it's all about?

OpenStudy (amistre64):

W = k/(d^2) 130 = k/(3978)^2 k = 130((3978)^2) k = 2,057,182,920 ------------------------- W = 2,057,182,920 // d^2 W = 2,057,182,920 // (3978+173)^2 W = 2,057,182,920 // 17,230,801 W = 119.38986 lbs so I would put in... 119.39 as an answer if not, then 119.38

OpenStudy (amistre64):

as a kid, it was either the hokey pokey are duck duck goose :)

OpenStudy (amistre64):

rock paper scissors hadnt been invented yet lol

OpenStudy (anonymous):

lol

OpenStudy (anonymous):

She says youre awesome!

OpenStudy (amistre64):

now now... she only says it cause its true :)

OpenStudy (amistre64):

or shes having a stroke....have her repeat the alphabet just to be sure ;)

OpenStudy (anonymous):

She couldnt do that anyways!

OpenStudy (anonymous):

We only have like two weeks left.........

OpenStudy (amistre64):

......cancer?

OpenStudy (anonymous):

almost.........math class.

OpenStudy (anonymous):

I have one more............ o........wait......

OpenStudy (anonymous):

Well, thanks!

OpenStudy (amistre64):

type in a new post ;) that way more people can have a chance to help.... they are jealous of my fan count I just know it ;)